Break-Even Sales and Sales Mix for a Service Company

Zero Turbulence Airline provides air transportation services between Los Angeles, California; and Kona, Hawaii. A single Los Angeles to Kona round-trip flight has the following operating statistics:

Fuel $6,956
Flight crew salaries 5,328
Airplane depreciation 2,516
Variable cost per passenger—business class 45
Variable cost per passenger—economy class 35
Round-trip ticket price—business class 525
Round-trip ticket price—economy class 285

It is assumed that the fuel, crew salaries, and airplane depreciation are fixed, regardless of the number of seats sold for the round-trip flight. If required round the answers to nearest whole number.

a. Compute the break-even number of seats sold on a single round-trip flight for the overall product, E. Assume that the overall product is 20% business class and 80% economy class seats.

Total number of seats at break-even seats

b. How many business class and economy class seats would be sold at the break-even point?

Business class seats at break-even seats
Economy class seats at break-even seats
Answer #1

a)

contribution margin of business class ticket = (525-45) = 480
contribution margin of economy class ticket = (285-35) = 250

weighted contribution margin per unit = 20%*480+80%*250 = 296

Break even number of seats = fixed cost/weighted contribution margin per unit

=(6956+5328+2516)/296 = 50

b.

Business class eats at break-even = 50*20% = 10 seats
Economy class seats at break-even = 50*80% = 40 seats